OpenStudy (anonymous):

helpp (1/2 radical y) (-6radical y) radical9a*radical 9b

OpenStudy (anonymous):

\[(1/2\sqrt{y}) (-6\sqrt{y)}\]

OpenStudy (anonymous):

\[\sqrt{9a} * \sqrt{9b}\]